### v2

# An LLM-Infused Tool for Workers’ Collective Intelligence

Abstract.
Collective intelligence among gig workers yields considerable advantages, including improved information exchange, deeper social bonds, and stronger advocacy for better labor conditions. Especially as it enables workers to collaboratively pinpoint shared challenges and devise optimal strategies for addressing these issues. However, enabling collective intelligence remains challenging, as existing tools often overestimate gig workers’ available time and uniformity in analytical reasoning. To overcome this, we introduce GigSense, a tool that leverages large language models alongside theories of collective intelligence and sensemaking. GigSense enables gig workers to rapidly understand and address shared challenges effectively, irrespective of their diverse backgrounds. GigSense not only empowers gig workers but also opens up new possibilities for supporting workers more broadly, demonstrating the potential of large language model interfaces to enhance collective intelligence efforts in the evolving workplace.
Gig work, Crowdsourcing, Future of work, Generative AI, Human-AI Collaborations
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computingccs: Collaborative and social computing systems and tools
Figure 1. Overview of our process for designing and studying a generative AI tool for gig workers.
1. Introduction
Harnessing collective intelligence among gig workers can significantly enhance their ability to improve labor conditions (Chen et al., 2023; Qadri, 2023; Grohmann et al., 2023; Piven and Cloward, 2012). This collaborative approach enables workers to collectively deliberate on their challenges, devise strategic solutions, and collaboratively implement their action plans to tackle these issues effectively(Bessa et al., 2022; Dubal, 2017).
However, despite occasional successes, collective intelligence among gig workers is uncommon, leaving many issues unresolved (Bessa et al., 2022). This scarcity largely stems from the lack of technologies designed for facilitating collective problem-solving among gig workers (Johnston, 2020; Shaw et al., 2014; Zhang et al., 2014; Gray and Suri, 2019; Lehdonvirta, 2022). Platforms like Dynamo and Coworker.org, which allow for sharing, prioritizing, and solving issues through upvote-based lists, fail to provide a comprehensive view of problems and solutions (Salehi et al., 2015; DeMaria, 2016). Such list-based interfaces limit the exploration of issues from multiple angles and understanding the full scope of potential solutions (Li et al., 2023; Khovanskaya et al., 2019; Calacci, 2022; Mattsson et al., 2015; Peter and Card, 2005; Weick, 1995), often leading workers to prioritize less critical concerns (Wood and Lehdonvirta, 2019). This impedes meaningful progress in addressing the fundamental challenges of gig work, especially during the initial phases of collective intelligence focused on recognizing shared problems and formulating solutions (Schwellnus et al., 2019).
Another hurdle is that the diverse and significant time commitments of gig workers may limit their ability to contribute to collective intelligence, underscoring the importance of developing inclusive tools that can accommodate workers’ varying schedules, reducing the barriers to collective intelligence and enabling easier participation in collaboratively tackling challenges (Packham, 2008; Pantea, 2013). To address these challenges, we introduce GigSense, a novel platform for fostering collective intelligence among gig workers by enabling collaborative problem-solving. By integrating Sense-Making Theory and leveraging the capabilities of large-language models (LLMs), GigSense features an LLM-enhanced interface that allows workers to deeply analyze their challenges and solutions from multiple perspectives (Peter and Card, 2005). Unlike existing platforms that offer a simple list-based view of issues (Salehi et al., 2015), GigSense enables a detailed exploration of problems, offering both a close-up and a broad overview to understand workplace dynamics better. Additionally, GigSense uses LLMs to facilitate collective brainstorming, helping workers develop solutions together. Fig 1 presents an overview of GigSense. As a significant advancement in fostering collective intelligence among gig workers, GigSense addresses the unique challenges of this workforce, which typically operates in isolation without a common platform for problem identification and solution generation (Calacci, 2022; Yao et al., 2021). GigSense bridges this gap by offering a communal space for gig workers to collectively understand and address their issue. Leveraging shared digital resources to build community (Sawyer et al., 2014; Barley, 1986), GigSense emerges as a vital tool for a fragmented workforce, potentially catalyzing the formation of supportive gig worker communities (Tajfel et al., 1979; Cohen et al., 2003; Van Zomeren et al., 2008; Tajfel and Turner, 2004). Beyond organization, it could reveal systemic issues, encouraging workers to see problems like client disputes and payment delays as collective challenges. This recognition can promote solidarity, crucial for collective intelligence, and empower workers to develop effective negotiation strategies, potentially improving their working conditions.
In this paper, we highlight a system that integrates LLMs with an interactive interface to support gig workers in problem-solving and kickstart collective intelligence among workers.
2. GigSense
GigSense is an AI-enhanced platform designed to assist gig workers in the early phases of sensemaking and collective intelligence, which includes problem identification and solution proposal (Shaw et al., 2014). Our platform thus focuses on aiding workers in their sensemaking process (identifying collective issues and potential solutions). For this purpose, GigSense incorporates Pirolli and Card’s sensemaking theories and collective intelligence theories into various modules within its system (Shaw et al., 2014; Weick, 1995; Mulgan, 2018a). These modules are designed to automate the sensemaking process, which unfolds seamlessly within the entirety of GigSense’s interface, with workers playing a crucial role in steering this process.
GigSense’s modules are based on previous designs (Flores-Saviaga et al., 2022; Li et al., 2018), but uniquely integrate human-AI collaboration to address specific needs and contexts of gig workers. GigSense recognizes the time constraints faced by gig workers when engaging in collective intelligence and hence provides interfaces that facilitate rapid sensemaking of problems and proposals of solutions. For this purpose, it utilizes Large Language Models (LLMs) to provide concise summaries of issues found on gig platforms, enabling workers to quickly understand prevalent problems. By leveraging the OpenAI GPT-4 API (OpenAI, 2024), it categorizes data and generates these summaries, aiding workers in gaining insights and brainstorming solutions. Our system also offers an interactive interface, allowing users to “zoom in” and “zoom out” and explore specific discussions related to a problem. While LLM supports the creation of problem summaries, GigSense’s interface design empowers workers to independently assess the information. It also provides a collaborative workspace for workers to propose solutions to identified problems. It uses LLMs to inspire workers with potential solutions but prioritizes human-generated ideas over AI suggestions to emphasize the value of human input. Notice that GigSense’s design, focusing on quick sensemaking, contrasts with traditional extended analysis but offers notable advantages. It enables time-constrained individuals like gig workers with multiple jobs, parents, and students to participate, enhancing diversity and inclusivity. This approach makes sensemaking more accessible, as it requires less long-term commitment (Etz et al., 2018). A faster process encourages broader participation and engagement, as people are more likely to contribute when their time investment is minimal yet impactful (Sandberg and Tsoukas, 2015). This approach can be especially beneficial for future engagements around collective intelligence, such as implementing solutions, where sustained participant involvement is crucial (Etz et al., 2018). Additionally, rapid sensemaking can foster concise and focused information processing, reducing data overload and maintaining participant interest, which can wane in longer processes (Mulgan, 2018b; Sandberg and Tsoukas, 2015). Next, we describe each GigSense module in detail:
Data Gathering Module.
Gig workers supply Gigsense with a roster of subreddits from which they intend to pinpoint potential problems and datasets containing assessments for gig work platforms (This is the “Step: Search and Filter” in Pirolli et al.’s sensemaking loop (Peter and Card, 2005)). Next, Gigsense connects to the Reddit API to read and extract all the posts from the subreddits that gig workers initially provide. GigSense additionally uses a web scrapper to extract data from reviews left on Apple’s and Google’s app stores by gig workers. Note that our data gathering module only collects reviews that have between one and three-star ratings. The module considers that these review data would represent complaints and problems that gig workers are experiencing. Gigsense also lets workers manually enter issues into the system if they choose to do so (“Step: Read and Extract” in the sensemaking loop). Using the real-world gig workers’ complaint datasets (actual gig workers’ subreddits and complaints) in our system design aims to bring inclusiveness about gig worker concerns and complaints. The show button allows workers to see all the data sources used in GigSense.
Problem Summary Module.
Acknowledging the potential enormity of the data gathered through the Data Gathering Module and its potential complexity for human interpretation, this module centers its efforts on summarizing the data (“Step: Schematize” in the sensemaking loop). For this purpose, Gigsense uses LLMs to facilitate the exploration and sensemaking of the extensive datasets of complaints, enabling workers to categorize and summarize the large datasets into problem categories. This module also has two buttons to allow workers to navigate to the Data Visualization module, which offers a “zoom-in” view for meticulous examination of the data or a “zoom-out” view for a panoramic grasp of their issues.
Data Viz Module
This module enables workers to analyze problems at various levels, aiding their sensemaking process. It utilizes semantic zooming, a method for visualizing data across different abstraction levels (Bederson and Hollan, 1994; Suh et al., 2023a, b). Workers have access to two views showcasing varying detail levels for problem analysis. The zoomed-out view of this module presents an interactive chart of the problems (“Step: Schematize” in the sensemaking loop) where workers can grasp a high-level understanding of the specific types of problems (e.g. Payment, Platform Policy, Scam, Customer Support, Usability, etc) faced by them. Gigsense utilizes LLMs to streamline the sensemaking process by categorizing dataset posts into distinct categories. These categorized posts are later used by the system to create the different graphs and the corresponding category descriptions. The use of LLMs speeds up data organization, allowing workers to engage in sensemaking more efficiently. The interactive chart also serves as a powerful tool for raising awareness among gig workers about their shared challenges. For instance, workers can engage with these graphs by hovering over them, enabling access to the exact number posts (complaints) within each category. This functionality contributes to the worker’s sensemaking process by facilitating the easy identification of prevalent issues.
The volume of information available to workers shifts according to the zoom level. To let the worker delve deeper into individual problems, Gigsense offers a zoom-in perspective. Choosing the zoom-in view shifts the display from graphs to the full text of posts relevant to the problem being analyzed. This view allows workers to view the nuanced aspects of issues that pertain to the topic analyzed and upvote on individual problems. By visually displaying the prevalence of issues (number of upvotes), it creates a sense of solidarity (Savage and Jarrahi, 2020). Workers can see that they are not alone in their struggles, fostering a collective identity and shared purpose (Lee et al., 2015). It provides a static depiction of the number of workers grappling with problems across different gig work platforms, acting as a catalyst in building communities of gig workers facing similar challenges to encourage collective intelligence. The zoomable interface facilitates sensemaking by allowing users to toggle between high-level context and detailed analysis. Note that as (Mulgan, 2018b) discusses, collective intelligence requires focused attention on what matters. The system is thus designed to allow workers to analyze problems from detailed and broader perspectives.
Collaborative Solution Module.
This module further facilitates the sensemaking process and focuses on helping gig workers to develop concrete solutions to address the problem analyzed (“Step: Build Case’ ” in the sensemaking loop). It incorporates sub-modules such as the: “Sensemaking Chat”, “Shared Document” and “Collaborative Solution Space”. The Sensemaking Chat submodule allows workers to engage in conversations to discuss and investigate the problems they encounter in their work. They can communicate through asynchronous text messages to accommodate different schedules. The “Shared Document” enables workers to understand the existing problems and collaboratively create action plans (solutions) to address them. Gigsense also includes a functionality that allows users to make annotations in the shared document. This way, workers can collectively review and approve their proposed solutions. Finally, the “Collaborative Solution Space”, just like the sensemaking process, features a space where workers can showcase the final solution they mutually agreed upon (Weick, 1995; Dervin, 1983) (“Step: Tell Story” in the sense-making loop). This module supports collaborative work that initiates early stages of collective intelligence(Shaw et al., 2014).
AI-Enhanced Solution Module.
It is important to acknowledge that for some workers it may be hard to propose solutions (Snook and Olsen, 2006; Packham, 2008). To address this, the module leverages LLMs to offer workers suggestions on potential solutions and concrete collaboration plans, providing inspiration and initial guidance. However, given our values of prioritizing human connections among workers, AI-generated solutions are presented with lower priority in GigSense’s interface. GigSense emphasizes that the AI-generated solution is not the definitive solution and explicitly states that this solution was generated using Generative AI. Note that in our system, AI-generated solutions are not personalized for individual users. However, the prompts sent to the LLM include all data related to the workers’ specific problem. Using real-world gig worker data likely enhances our system’s ability to provide contextually relevant AI solutions. Our appendix provides information about all the prompts we utilized for the LLM. Note also that GigSense uses pre-defined prompts to help gig workers effectively interact with LLMs, without requiring specialized skills (Zamfirescu-Pereira et al., 2023). The pre-defined prompts are also used because GigSense is not designed to address all queries, but instead, aims to offer concrete solutions for problem-solving inspiration. Furthermore, workers are informed about the possibility of errors in AI-generated solutions via a disclaimer. This helps not only to ensure that solutions generated by humans receive priority, but also foster more transparent and responsible use of Generative AI while still utilizing AI to assist and inspire workers in their problem-solving endeavors. This module is designed to enhance the “Step: Build Case” and “Step: Tell Story” stages of the sensemaking loop. It achieves this by promoting a collaborative process between humans and AI, fostering co-creation and mutual contribution to the development of solutions.
3. Conclusion
This paper introduces GigSense, an AI-powered tool designed to assist gig workers in collectively addressing their challenges. GigSense enables quick sensemaking, promoting effective problem-solving among workers by breaking down barriers to collective intelligence.
Acknowledgments This work was partially supported by an NSF grant FW-HTF-19541.